OpenStudy (anonymous):

6.78j-5.2=4.33j+2.15

OpenStudy (texaschic101):

we can work this problem with the decimal in there, or we can work it without the decimals. Which do you prefer ?

OpenStudy (anonymous):

with decimal

OpenStudy (texaschic101):

6.78j - 5.2 = 4.33j + 2.15 -- subtract 4.33j from both sides 6.78j - 4.33j - 5.2 = 2.15 -- add 5.2 to each side 6.78j - 4.33j = 2.15 + 5.2 -- combine like terms 2.45j = 7.35 -- divide both sides by 2.45 j = 7.35/2.45 j = 3 any questions ?

OpenStudy (anonymous):

nope and thanks soo much

OpenStudy (texaschic101):

your quite welcome :)